# VMem: Consistent Interactive Video Scene Generation with Surfel-Indexed View Memory

### v2

@@ Abstract
We propose a novel memory mechanism to build video generators that can explore environments interactively. Similar results have previously been achieved by out-painting 2D views of the scene while incrementally reconstructing its 3D geometry, which quickly accumulates errors, or by video generators with a short context window, which struggle to maintain scene coherence over the long term. To address these limitations, we introduce Surfel-Indexed View Memory (VMem), a mechanism that remembers past views by indexing them geometrically based on the 3D surface elements (surfels) they have observed. VMem enables the efficient retrieval of the most relevant past views when generating new ones. By focusing only on these relevant views, our method produces consistent explorations of imagined environments at a fraction of the computational cost of using all past views as context. We evaluate our approach on challenging long-term scene synthesis benchmarks and demonstrate superior performance compared to existing methods in maintaining scene coherence and camera control.
Figure 1: VMem enables interactive scene generation from a single image along user-specified trajectories in an auto-regressive manner. The green region shows results with memory, maintaining consistency when generating previously observed parts of the scene. The red region, without memory, exhibits visual drift highlighted with red ellipses, demonstrating the effectiveness of VMem for consistent scene generation.
@@ 1 Introduction
We consider the problem of generating long videos that explore an imagined space following a camera path specified interactively by the user. In this paradigm, the user tells the model which camera path to follow for the next few frames, observes the generated content, and then decides where to explore next based on what they have seen. For example, a possible exploration of a house may involve visiting the kitchen, the living room, and the bathroom, eventually returning to the kitchen. Throughout the video, the scene must remain consistent, ensuring that the kitchen looks the same upon return. The ability to generate such videos is essential for immersive applications such as games, where players can navigate generated worlds. However, even recent large-scale interactive video generators such as Google’s Genie 2 [21] struggle to achieve this goal.
@@
This problem has so far been addressed by two types of methods. First, out-painting-based methods [37, 7, 12, 19, 22, 42, 27, 16, 17, 25] iterate between generating new 2D views of the scene and estimating its 3D geometry. They use the estimated geometry to partially render a new viewpoint and then employ an out-painting model to complete the missing parts, thus adding one more image to the collection. However, errors in out-painting, 3D reconstruction, and stitching accumulate over time, leading to severe degradation of the generated content after a short while.
@@
Second, multi-view/video-based methods [26, 41, 23, 36, 21] condition the generation of a novel view on previously generated views using a geometry-free approach that does not explicitly estimate the scene geometry. While this avoids the accumulation of errors in the reconstructed 3D scene geometry, it comes at a high computational cost, limiting the number of conditioning views to a small context window consisting of only the most recent few frames. This constraint negatively impacts the long-term consistency of the generated images.
@@
Figure 2: Method overview. Given novel camera viewpoints $\{\mathbf{c}_{T+m}\}_{m=1}^{M}$, we first use our proposed memory module Surfel-Indexed View Memory $\mathcal{V}^{(s)}$ to retrieve the most relevant $K$ past views $v_{t}\in\mathcal{V}^{*}\subset\mathcal{V}^{(s)}$ as conditioning reference views. For each retrieved view, the camera parameters $\mathbf{c}_{t}$ are represented using Plücker embeddings, which are spatially resized via interpolation to match the resolution of the encoded image latents. The corresponding reference images $\mathbf{x}_{t}$ are encoded using a VAE to obtain latent feature maps. These resized Plücker embeddings and VAE-encoded latents are then fed into the camera-conditioned image-set generator $\psi$ to synthesize the novel views $\{\mathbf{x}_{T+m}\}_{m=1}^{M}$. Once the new views are generated, we update the surfel-indexed memory $\mathcal{S}^{(s)}\rightarrow\mathcal{S}^{(s+1)}$ by appending the new view indices $\{T+m\}_{m=1}^{M}$ to existing surfels or creating new ones based on the estimated geometry of the generated views. This process is repeated autoregressively for each subsequent generation step.
@@
In this work, we revisit the second class of methods and propose conditioning not on the most recently generated views, but on the most relevant ones for generating the next image, thereby maintaining a high degree of consistency within a limited computational budget. Given a novel viewpoint of the scene, the most relevant past views are those that have already observed the parts of the scene currently being generated. This implies that, for each part of the scene, we must remember which views have previously observed it and allow one to retrieve them from memory.
@@
To achieve this, we introduce the Surfel-Indexed View Memory, abbreviated as VMem, a memory mechanism that anchors previous views to the 3D surface elements they observe. Given a new viewpoint, we retrieve the past views that best capture the currently observed surfaces and use them to condition the generation of the novel view. To create the memory, we estimate the geometry of each new view using an off-the-shelf point map predictor. This is similar to out-painting methods but, crucially, we do not use the estimated geometry as the final representation of the scene; instead, we use it to construct VMem, which is a memory of past views. We represent the scene geometry using surfels which, compared to meshes, are more robust, and, compared to point clouds, can represent occlusions. Each surfel stores in its attributes a set of indexes corresponding to past viewpoints that observed it.
@@
To retrieve relevant past viewpoints, we render the surfels with their attributes from the novel viewpoint and splat them onto an image grid. Each pixel in the resulting image then corresponds to a set of viewpoint indexes. We select the top $n$ most frequently represented viewpoint indexes and use them to retrieve past views from the database, where each view is represented by an RGB image and its corresponding camera parameters. A key advantage of our approach over inpainting-based methods is that it does not require highly accurate scene geometry. As long as we successfully retrieve the most relevant past views, our method remains robust.
@@
    We introduce Surfel-Indexed View Memory, a plug-and-play module to index past views geometrically and use them to condition the generation of novel views.
@@
    We show that our method can generate long-term coherent videos of scenes, outperform existing approaches.
@@
    We validate our approach on a number of challenging benchmarks, outperforming the current open-source state-of-the-art methods.
@@ 2 Related Work
Novel view synthesis (NVS) is a challenging and ill-posed problem. NVS methods can generally be categorized into two main approaches: view interpolation methods [20, 3, 4, 5, 6, 38, 31, 9], which generate views that stay close to the given inputs, and view extrapolation or perpetual view generation, where novel views extend significantly beyond the original scene, often introducing large amounts of new content. The latter is particularly difficult when only a single image is available. Between these two categories lie methods focusing on single-view scene reconstruction, which have been mostly successful for scenes containing a single object [18, 28, 13, 30, 1, 39] or highly bounded scenes [29]. However, their extrapolation capabilities remain very limited. The most relevant to our work are models for single-image perpetual view generation, which typically fall into two main categories: those that incorporate explicit geometric modeling with inpainting techniques and those directly based on image or video generation.
Inpainting-based perpetual view generation.
@@
Inpainting-based methods [37, 7, 12, 19, 22, 42, 27, 16, 17, 25] use pre-trained 3D reconstruction methods to generate explicit 3D representations—such as meshes, point clouds, or Gaussian splats—from an initial image. These representations are reprojected into novel views, where 2D inpainting models fill in missing regions. For example, SceneScape [7] reconstructs a mesh from the first image using a pre-trained depth estimator. A diffusion-based inpainting model then completes the projected novel view, which is reprojected back to refine and extend the mesh. This process is iterated to generate additional views. MultiDiff [19] also relies on a depth estimator to warp the reference image into multiple novel views, training a diffusion model to inpaint missing regions simultaneously. CamCtrl3D [22] follows a similar approach but incorporates a ray map to better condition the diffusion model for view synthesis. ViewCrafter [42] uses a pre-trained pointmap estimator to create a 3D point cloud and trains a video diffusion model for inpainting. GenWarp [27], uniquely, warps a 2D coordinate representation of the reference view into novel views and uses it to condition generation.
@@
While these methods improve geometric consistency, they are highly susceptible to errors in depth or pointmap estimation, which can propagate distortions across generated views. Once the 3D representation is constructed, any inaccuracies from depth or pointmap estimation become difficult to correct. Moreover, scaling these approaches to large scenes is computationally intensive, as storing and processing explicit high-fidelity 3D representations requires substantial memory.
@@
Figure 3: Visualization of our surfel-based memory index. Each surfel in the 3D scene stores indices of all views that have observed it. For visualization, we color-code each surfel by the indices of the contributing views. This spatial index enables efficient retrieval of relevant past views: when generating a new view, we identify which surfels are visible from the target viewpoint and retrieve the views that previously observed those same regions, naturally accounting for occlusion.
Multi-view-based perpetual view generation.
Another category of methods [26, 23, 32, 41, 36, 8, 46] avoids explicitly modeling the 3D geometry of the scene. Instead, these approaches generate novel views by conditioning on previously rendered or generated views.
@@
GeoGPT [26] establishes this direction by using an autoregressive likelihood model to synthesize novel perspectives from a single input image. Building upon this, LookOut [23] improves consistency by generating view sequences along a predefined camera trajectory, conditioning on up to two preceding views. It also incorporates a post-processing step that reuses previously generated frames as additional context for future predictions. Other methods, such as [32, 41], enhance novel view synthesis by employing cross-view attention or enforcing epipolar constraints within diffusion models. MotionCtrl [36], in particular, leverages geometric priors from a pre-trained video diffusion model to synthesize camera-controllable videos. All of the aforementioned methods are trajectory-based, requiring the input views to follow a continuous spatial path. In contrast, image-set models like CAT3D [8] are capable of synthesizing novel views from a sparse and unordered set of input images, without any trajectory constraint. Most recently, SEVA [46] has emerged as a strong open-source alternative in this image-set modeling paradigm.
However, all these methods are computationally expensive, with an $O(n^{2})$ complexity for attention, which restricts the number of conditioning views to a small context window containing only a few of the most recently generated frames. This limitation negatively impacts the long-term consistency of the generated images.
@@
Several concurrent works have explored memory-based approaches for consistent scene generation. StarGen [44] and WorldMem [40] use distance and field-of-view-based spatial memory retrieval, which can recover spatially correlated views but lack geometric reasoning for occlusion handling. Genie [2, 21] maintains memory through recurrent state features. Gen3C [24] conditions on stored point clouds similar to ViewCrafter [43], inheriting the limitations of inpainting-based methods. In contrast, our surfel-based memory explicitly models occlusion and provides more principled view selection.
@@ 3 Method
Let $\mathbf{x}_{1}\in\mathbb{R}^{H\times W\times 3}$ be an RGB image of a scene that we wish to explore, and let $\{\mathbf{c}_{t}\}_{t=1}^{T}$ be a sequence of camera parameters specifying a path through the scene. Our goal is to generate a corresponding video, i.e., a sequence of images $\{\mathbf{x}_{t}\}_{t=1}^{T}$, where $\mathbf{x}_{1}$ is the given input image, and each subsequent frame $\mathbf{x}_{t}$ is consistent with the previous ones and reflects the specified camera $\mathbf{c}_{t}$. Moreover, we aim to do so autoregressively, generating $M$ novel views at a time for the given cameras (with $M$ kept small to ensure interactivity) while ensuring consistency with previously generated views. This enables interactive scene exploration, where the user can specify the next camera positions at each step.
We define a view as a tuple $v_{t}=(\mathbf{x}_{t},\mathbf{c}_{t})$ consisting of an image $\mathbf{x}_{t}$ and its corresponding camera $\mathbf{c}_{t}$ at timestep $t$. At generation step $s$, we have generated $T=sM,$ $s\geq 0$ frames so far. The goal is to generate $M$ new RGB images $\{\mathbf{x}_{T+m}\}_{m=1}^{M}$ for the next target camera parameters $\{\mathbf{c}_{T+m}\}_{m=1}^{M}$, given the previously generated views $\mathcal{V}^{(s)}=\{v_{1},v_{2},\dots,v_{T}\}$. After generation, we update $T\leftarrow T+M$.
@@ 3.1 Surfel-Indexed View Memory
Consider the problem of generating the next views $\{\mathbf{x}_{T+m}\}_{m=1}^{M}$ of the scene as seen from the cameras $\{\mathbf{c}_{T+m}\}_{m=1}^{M}$. The generator must consider the previously generated views and ensure $\{\mathbf{x}_{T+m}\}_{m=1}^{M}$ are consistent with those, while generating new content as needed. However, not all past views are equally correlated with the novel views $\{\mathbf{x}_{T+m}\}_{m=1}^{M}$. We prioritize past views that have likely observed the largest portion of the scene currently being generated. For example, if the current locale is surrounded by walls, parts of the scene behind those walls are likely less relevant for generating a new view of that locale. Conversely, views that share a similar visible region with the target view can provide more useful information for generating its content.
This principle drives our design for a view memory mechanism. We maintain a coarse model of the scene geometry using surfels (visualized in Fig. 3), simple surface primitives that account for occlusion. Each surfel stores the indices of the past views that observed it. To generate a new view, the surfels visible from the new viewpoint are retrieved, using the associated indices to vote for which views to consider for generation. This process is illustrated in Fig. 4.
@@
Figure 4: Surfel-Indexed View Memory. Reading from the memory involves rendering the surfels $\mathcal{S}^{(s)}$ along with their attributes, which contain past view indices represented by frame indices. We then select the $K$ most frequently represented frame indices in the rendered image to retrieve the most relevant past views from $\mathcal{V}^{(s)}$. Writing to the memory involves estimating the geometry of the newly generated views $\{\mathbf{x}_{T+m}\}_{m=1}^{M}$ in the form of surfels and merging them with the existing surfels. The frame indices $\{T+m\}_{m=1}^{M}$ are appended to the surfels observed in these views, and the novel views are stored in memory, updating $\mathcal{V}^{(s)}\rightarrow\mathcal{V}^{(s+1)}$ and $\mathcal{S}^{(s)}\rightarrow\mathcal{S}^{(s+1)}$.
@@
To retrieve the most relevant past views from the memory $\mathcal{V}^{(s)}$ for the novel cameras $\{\mathbf{c}_{T+m}\}_{m=1}^{M}$, we first compute their average pose $\bar{\mathbf{c}}_{s}\in\mathrm{SE}(3)$ as a reference (refer to Appendix A). Then we render the surfels $\mathcal{S}^{(s)}$ from the averaged camera pose $\bar{\mathbf{c}}_{s}$. Each surfel is rendered as a splat with its attributes, accounting for relative depth and occlusions, and contributing to the image based on its coverage. The rendered attributes correspond to the indices of the previous views. The core intuition is that views observing the largest portion of the scene from the perspective of the novel camera are most relevant for novel view synthesis. To identify these views, we rank past view indices based on their frequency across all rendered pixels.
We then select the top-$K$ most frequently represented indices $\mathcal{I}^{*}$ and use them to retrieve the corresponding past views, $\mathcal{V}^{*}=\{v_{t}\}_{t\in\mathcal{I}^{*}}$, from the memory $\mathcal{V}^{(s)}$. This subset of views is used to feed the new view generator, as described in Sec. 3.2.
To avoid oversampling repeatedly visited regions, we apply a non-maximum suppression algorithm that reduces redundancy in memory and promotes broader scene coverage among the top-$K$ views. During retrieval, we retain only the most frequently referenced view among those with similar poses. Additionally, during writing to the memory, we merge surfels by comparing their associated camera poses—if two poses are highly similar, we discard the older one.
@@
After generating the new views $\{v_{T+m}=(\mathbf{x}_{T+m},\mathbf{c}_{T+m})\}_{m=1}^{M}$, we add them to the memory $\mathcal{V}^{(s)}$ and update the surfel-based memory index $\mathcal{S}^{(s)}$. We use an off-the-shelf point map estimator $\phi$, such as CUT3R [33], to output their corresponding point maps $\{\mathbf{P}^{*}_{T+m}\}_{m=1}^{M}$. Specifically, we jointly estimate the point maps for the newly generated views along with the retrieved past views $\mathcal{V}^{*}$. This joint estimation ensures that the new point maps are aligned with the coordinate frame of the existing scene geometry in the surfel-based memory index $\mathcal{S}^{(s)}$.
@@
$\mathbf{n}_{k^{\prime}}=\frac{(\mathbf{p}_{u+1,v,t}-\mathbf{p}_{u-1,v,t})\times(\mathbf{p}_{u,v+1,t}-\mathbf{p}_{u,v-1,t})}{\left\|(\mathbf{p}_{u+1,v,t}-\mathbf{p}_{u-1,v,t})\times(\mathbf{p}_{u,v+1,t}-\mathbf{p}_{u,v-1,t})\right\|}.$

### v3

@@ Abstract
We propose a novel memory module for building video generators capable of interactively exploring environments. Previous approaches have achieved similar results either by out-painting 2D views of a scene while incrementally reconstructing its 3D geometry—which quickly accumulates errors—or by using video generators with a short context window, which struggle to maintain scene coherence over the long term. To address these limitations, we introduce Surfel-Indexed View Memory (VMem), a memory module that remembers past views by indexing them geometrically based on the 3D surface elements (surfels) they have observed. VMem enables efficient retrieval of the most relevant past views when generating new ones. By focusing only on these relevant views, our method produces consistent explorations of imagined environments at a fraction of the computational cost required to use all past views as context. We evaluate our approach on challenging long-term scene synthesis benchmarks and demonstrate superior performance compared to existing methods in maintaining scene coherence and camera control.
Figure 1: VMem enables autoregressive scene generation from a single image along user-defined trajectories. The green region shows results with the proposed memory module, maintaining coherence when generating previously seen parts of the scene. The red region, without memory, exhibits degradation highlighted with red ellipses, demonstrating VMem is effective for consistent scene generation.
@@ 1 Introduction
We consider the problem of generating long videos that explore an imagined space following a camera path specified interactively by the user. In this paradigm, the user tells the model which camera path to follow for the next few frames, observes the generated content, and then decides where to explore next based on what they have seen. For example, exploring a house may involve visiting the kitchen, the living room, and the bathroom, eventually returning to the kitchen. Throughout the video, the scene must remain consistent, ensuring that the kitchen looks the same upon return. Generating such videos is essential for immersive applications such as games, where players can navigate generated worlds. However, even recent large-scale interactive video generators such as Google’s Genie 2 [20] struggle to achieve this goal.
@@
This problem has so far been addressed by two types of methods. First, outpainting-based methods [36, 7, 11, 18, 21, 43, 26, 15, 16, 24] iterate between generating new 2D views of the scene and estimating its 3D geometry. They use the estimated geometry to partially render a new viewpoint and then employ an outpainting model to complete the missing parts, thus adding one more image to the collection. However, errors in outpainting, 3D reconstruction, and stitching accumulate over time, leading to severe degradation of the generated content after a short while.
@@
Second, multi-view/video-based methods [25, 42, 22, 35, 20] condition novel view generation on previous views using a geometry-free approach that does not explicitly estimate the scene geometry. While this avoids the accumulation of errors in the reconstructed 3D scene geometry, it comes at a high computational cost, limiting the number of conditioning views to a small context window of recent frames. This constraint hurts the long-term consistency of the generated images.
@@
Figure 2: Method. Given target camera viewpoints $\{\mathbf{c}_{T+m}\}_{m=1}^{M}$, we query our Surfel-Indexed View Memory to retrieve the most relevant $K$ past views $\mathcal{V}^{*}\subset\mathcal{V}^{(s)}$ where $\mathcal{V}^{*}=\{v_{t}\}_{t=1}^{K}$ as references. Retrieved reference images $\mathbf{x}_{t}$ along with Plücker embeddings of both reference camera poses $\mathbf{c}_{t}$ and target camera poses $\{\mathbf{c}_{T+m}\}_{m=1}^{M}$ are fed into generator $\psi$ to synthesize novel views $\{\mathbf{x}_{T+m}\}_{m=1}^{M}$. After generation, the surfel-indexed memory is updated $\mathcal{S}^{(s)}\rightarrow\mathcal{S}^{(s+1)}$ by appending new view indices $\{T+m\}_{m=1}^{M}$ to existing surfels or creating new surfels based on geometry of the generated views. This is repeated autoregressively, enabling long-term consistent generation.
@@
In this work, we revisit the second class of methods and propose conditioning not on the most recently generated views, but on the most relevant ones for generating the next image, thereby maintaining a high degree of consistency within a limited computational budget. Given a novel viewpoint of the scene, the most relevant past views are those that have already observed the parts of the scene currently being generated. This implies that, for each part of the scene, we must remember which views have previously observed it and retrieve them from memory.
@@
To achieve this, we introduce Surfel-Indexed View Memory, abbreviated as VMem, a memory module that anchors previous views to the 3D surface elements they observe. Given a new viewpoint, we retrieve the past views that best capture the currently observed surfaces and use them to condition novel view generation. To create the memory, we estimate the geometry of each new view using an off-the-shelf point map predictor. This is similar to outpainting methods but, crucially, we do not use the estimated geometry as the final representation of the scene; instead, we use it to construct VMem, which is a memory of past views. We represent the scene geometry using surfels, which are more robust compared to meshes and can represent occlusions compared to point clouds. Each surfel stores in its attributes a set of indices corresponding to past viewpoints that observed it.
@@
To retrieve relevant past viewpoints, we render the surfels with their attributes from the novel viewpoint and splat them onto an image grid. Each pixel in the resulting image then corresponds to a set of viewpoint indices. We select the top $K$ most frequently represented viewpoint indices and use them to retrieve past views from the database, where each view is represented by an RGB image and its corresponding camera parameters. A key advantage of our approach over outpainting-based methods is that it does not require highly accurate scene geometry. As long as we successfully retrieve the most relevant past views, our method remains robust.
@@
    We introduce Surfel-Indexed View Memory, a plug-and-play module to index past views geometrically and use them to condition novel view generation.
@@
    We show that our method can generate long-term coherent videos of scenes and outperforms existing approaches.
@@
    We demonstrate that Surfel-Indexed View Memory achieves comparable performance with $4\times$ fewer context views, delivering a $12\times$ speedup.
  4. 4.
    We validate our approach on challenging benchmarks, outperforming the current open-source state-of-the-art methods.
@@ 2 Related work
Novel view synthesis (NVS) is a challenging and ill-posed problem that can be categorized into two main categories: view interpolation methods [19, 3, 4, 5, 6, 37, 30, 8], which generate views close to given inputs, and view extrapolation or autoregressive view generation, where novel views extend significantly beyond the original scene, introducing substantial new content. The latter is particularly difficult from single images. Between these lie single-view scene reconstruction methods, which succeed mainly for single-object scenes [17, 27, 12, 29, 1, 38] or highly bounded scenes [28]. However, their extrapolation capabilities remain limited. Most relevant to our work are single-image view extrapolation models, which fall into two categories: those incorporating explicit geometric modeling with inpainting, and those based directly on image or video generation.
Inpainting-based view extrapolation.
@@
Inpainting-based methods [36, 7, 11, 18, 21, 43, 26, 15, 16, 24, 40, 41] use pre-trained 3D reconstruction models to generate 3D representations—such as meshes, point clouds, or Gaussian splats—from images. These representations are reprojected into novel views, where 2D inpainting fills missing regions. SceneScape [7] reconstructs meshes from images using pre-trained depth estimators. Diffusion-based inpainting then completes the projected novel view, which is reprojected back to refine and extend the mesh. This process is iterated to generate novel views. MultiDiff [18] uses depth estimators to warp reference images into multiple novel views, training diffusion models to inpaint missing regions simultaneously. CamCtrl3D [21] follows a similar approach but adds ray maps to condition diffusion models for view synthesis. ViewCrafter [43] uses pre-trained pointmap estimators to create point clouds and trains video diffusion models for inpainting. GenWarp [26] warps 2D coordinate representations of reference views into novel views and uses them to condition generation.
@@
While these methods improve geometric consistency, they are susceptible to errors in depth or pointmap estimation, which propagate distortions across generated views. Once 3D representations are constructed, inaccuracies from depth or pointmap estimation become difficult to correct. Moreover, scaling to large scenes is computationally intensive, as storing and processing high-fidelity 3D representations requires substantial memory.
@@
Figure 3: Surfel-based memory index. Each surfel stores indices of views that observed it. We color-code each surfel by contributing view indices. This spatial index enables retrieval of relevant past views: when generating a novel view, we identify visible surfels from the target viewpoint and retrieve views that previously observed those same regions, naturally accounting for occlusion.
Multi-view-based view extrapolation.
@@
Another category [25, 22, 31, 42, 35, 8, 46] avoids explicitly modeling 3D scene geometry. Instead, these approaches generate novel views by conditioning on previously rendered or generated views. GeoGPT [25] uses autoregressive likelihood models to synthesize novel views from single images. Building on this, LookOut [22] improves consistency by generating view sequences along predefined trajectories, conditioning on up to two preceding views. Other methods [31, 42] enhance novel view synthesis by employing cross-view attention or enforcing epipolar constraints within diffusion models. MotionCtrl [35] leverages geometric priors from pre-trained video diffusion models to synthesize camera-controllable videos. These methods are trajectory-based, requiring input views to follow continuous spatial paths. In contrast, image-set models like CAT3D [8] and more recently SEVA [46] synthesize novel views from sparse and unordered input images, without trajectory constraints. However, these methods are computationally expensive, with $O(n^{2})$ complexity for attention, limiting conditioning views to small context windows containing only recently generated frames. This limitation impacts long-term consistency of generated images.
@@
Several concurrent works explore memory-based approaches for scene generation. Genie [2, 20] maintains memory through recurrent state features. Gen3C [23] conditions on stored point clouds like ViewCrafter [43], inheriting inpainting-based limitations. StarGen [44] and WorldMem [39] use distance- and field-of-view-based spatial memory retrieval, which can recover spatially correlated views but lack geometric reasoning for occlusions. In contrast, our surfel-based memory explicitly models occlusions and provides more principled view selection.
@@ 3 Method
Let $\mathbf{x}_{1}\in\mathbb{R}^{H\times W\times 3}$ be an RGB image of a scene that we wish to explore, and let $\{\mathbf{c}_{t}\}_{t=1}^{T}$ be a sequence of camera parameters specifying a path through the scene. Our goal is to generate a corresponding video, i.e., a sequence of images $\{\mathbf{x}_{t}\}_{t=1}^{T}$, where $\mathbf{x}_{1}$ is the given input image, and each subsequent frame $\mathbf{x}_{t}$ is consistent with the previous ones and reflects the specified cameras $\{\mathbf{c}_{t}\}_{t=1}^{T}$. Moreover, we aim to do so autoregressively, generating $M$ novel views at a time for the given cameras (with $M$ kept small to ensure interactivity) while ensuring consistency with previously generated views. This enables interactive scene exploration, where the user can specify the next camera positions at each step.
Specifically, we work with a sequence of views, where each view $v_{t}=(\mathbf{x}_{t},\mathbf{c}_{t})$ consists of an image $\mathbf{x}_{t}$ and its corresponding camera $\mathbf{c}_{t}$ at timestep $t$. At generation step $s$ ($s\geq 0$), we have generated $T=sM$ frames so far. The goal is to generate $M$ new RGB images $\{\mathbf{x}_{T+m}\}_{m=1}^{M}$ for the next target camera parameters $\{\mathbf{c}_{T+m}\}_{m=1}^{M}$, given the previously generated views $\mathcal{V}^{(s)}=\{v_{1},v_{2},\dots,v_{T}\}$. After generation, we update $T\leftarrow T+M$.
@@ 3.1 Surfel-indexed view memory
Consider the problem of generating the next views $\{\mathbf{x}_{T+m}\}_{m=1}^{M}$ of the scene as seen from the cameras $\{\mathbf{c}_{T+m}\}_{m=1}^{M}$. The generator must consider the previously generated views and ensure $\{\mathbf{x}_{T+m}\}_{m=1}^{M}$ are consistent with those, while generating new content as needed. However, not all past views are equally relevant to the novel views $\{\mathbf{x}_{T+m}\}_{m=1}^{M}$. We prioritize past views that have likely observed the largest portion of the scene currently being generated. For example, if the current locale is surrounded by walls, parts of the scene behind those walls are likely less relevant for generating a new view of that locale. Conversely, views that share a similar visible region with the target view can provide more useful information for generating its content.
This principle drives our design for a view memory module. We maintain a coarse model of the scene geometry using surfels (visualized in Fig. 3), simple surface primitives that account for occlusion. Each surfel stores the indices of the past views that observed it. To generate a new view, the surfels visible from the new viewpoint are retrieved, using the associated indices to vote for which views to consider for generation. This process is illustrated in Fig. 4.
@@
Figure 4: Surfel-Indexed View Memory. Reading procedure renders surfels $\mathcal{S}^{(s)}$ with their attributes, containing past view indices as frame indices. We then select the $K$ most frequent frame indices in the rendered image to retrieve relevant past views from $\mathcal{V}^{(s)}$. Writing procedure estimates geometry of newly generated views $\{\mathbf{x}_{T+m}\}_{m=1}^{M}$ as surfels and merges them with existing surfels. Frame indices $\{T+m\}_{m=1}^{M}$ are appended to surfels in these views, and novel views are stored, updating $\mathcal{V}^{(s)}\rightarrow\mathcal{V}^{(s+1)}$ and $\mathcal{S}^{(s)}\rightarrow\mathcal{S}^{(s+1)}$.
@@
To retrieve the most relevant past views from the memory $\mathcal{V}^{(s)}$ for the novel cameras $\{\mathbf{c}_{T+m}\}_{m=1}^{M}$, we first compute their average pose $\bar{\mathbf{c}}_{s}\in\mathrm{SE}(3)$ as a reference (refer to Appendix B). Then we render the surfels $\mathcal{S}^{(s)}$ from the averaged camera pose $\bar{\mathbf{c}}_{s}$. Each surfel is rendered as a splat with its attributes, accounting for relative depth and occlusions, and contributing to the image based on its coverage. The rendered attributes correspond to the indices of the past views. The core intuition is that views observing the largest portion of the scene from the perspective of the novel camera are most relevant for novel view synthesis. To identify these views, we rank past view indices based on their frequency across all rendered pixels.
We then select the top-$K$ most frequently represented indices $\mathcal{I}^{*}$ and use them to retrieve the corresponding past views, $\mathcal{V}^{*}=\{v_{t}\}_{t\in\mathcal{I}^{*}}$, from the memory $\mathcal{V}^{(s)}$. This subset of views serves as the context for the new view generator, as described in Sec. 3.2.
To avoid oversampling repeatedly visited regions, we apply a non-maximum suppression algorithm that reduces redundancy in memory and promotes broader scene coverage among the top-$K$ views. During retrieval, we retain only the most frequently referenced view among those with similar poses. During writing to the memory, we merge surfels by comparing their associated camera poses—if two poses are highly similar, we discard the older one.
@@
After generating the new views $\{v_{T+m}=(\mathbf{x}_{T+m},\mathbf{c}_{T+m})\}_{m=1}^{M}$, we add them to the memory $\mathcal{V}^{(s)}$ and update the surfel-based memory index $\mathcal{S}^{(s)}$. We use an off-the-shelf point map estimator $\phi$, such as CUT3R [32], to estimate their point maps $\{\mathbf{P}^{*}_{T+m}\}_{m=1}^{M}$. Specifically, we jointly estimate the point maps for the newly generated views along with the retrieved past views $\mathcal{V}^{*}$. This joint estimation ensures that the new point maps are aligned with the coordinate frame of the existing scene geometry in the surfel-based memory index $\mathcal{S}^{(s)}$.
@@
$\mathbf{n}_{k^{\prime}}=\frac{(\mathbf{p}_{u+1,v,t}-\mathbf{p}_{u-1,v,t})\times(\mathbf{p}_{u,v+1,t}-\mathbf{p}_{u,v-1,t})}{\left\|(\mathbf{p}_{u+1,v,t}-\mathbf{p}_{u-1,v,t})\times(\mathbf{p}_{u,v+1,t}-\mathbf{p}_{u,v-1,t})\right\|},$